# Black hole spectroscopy of collapsing and merging neutron stars

### v1

@@ Abstract
Black hole spectroscopy is an important pillar when studying gravitational waves from black holes and enables tests of general relativity. Most of the gravitational-wave signals observed over the last decade originate from binary black hole systems. Binary neutron star or black hole-neutron star systems are rarer but of particular interest for the next-generation ground-based gravitational-wave detectors. These events offer the exciting possibility of studying matter effects on the ringdown of “dirty black holes”. In this work, we ask the question: Does matter matter? Using numerical-relativity, we simulate a wide range of collapsing neutron stars producing matter environments, both in isolated scenarios and in binary mergers. Qualitatively, the resulting ringdown signals can be classified into “clean”, “modified”, and “distorted” cases, depending on the amount of matter that is present. We apply standard strategies for extracting quasinormal modes of clean signals, using both theory-agnostic and theory-specific assumptions. Even in the presence of matter, possible modifications of quasinormal modes seem to be dominated by ringdown modeling systematics. We find that incorporating multiple quasinormal modes allows one to drastically reduce mismatches and errors in estimating the final black hole mass at early times. If not treated carefully, deviations in the fundamental quasinormal mode might artificially be overestimated and falsely attributed to the presence of matter or violations of general relativity.
@@ I Introduction
By concentrating on the ringdown, it is possible to test the no-hair theorem’s assumptions. A BH formed from the dynamical processes involving compact objects “rings down” from an initial state of non-equilibrium and emits GW radiation in form of damped sinusoids. Pioneering works first targeted non-rotating BHs [24, 25, 26, 27], and then rotating ones [28]. They discovered how a BH responds to an incoming GW pulse: it oscillates with characteristic damped frequencies. The damping is due to GWs in the vicinity of the horizon falling into the BH, while GWs far away leave the system. The ringdown is a consequence of both conditions. This stage is anticipated to occur as a result of the coalescence of compact objects or in other instances where BH formation occurs, such as the collapse of a perturbed NS. Seminal studies probing the non-linear BH ringdown with numerical-relativity can be found in Ref. [29, 30, 31].
@@
In vacuum and in GR, the QNM spectrum is well known [34], while it requires strong simplifications to examine the QNM content in matter environments to make the equations decouple and make the analysis tractable. Maintaining the vacuum frequencies at a constant value from the start, allows for a simplified extraction of the complex-valued amplitudes $C_{\ell mn}$ from the signal. The amplitudes are fitted using standard least-squares fitting procedures [43, 44, 45], while the variable projection method [46, 47] is employed for the non-linear fitting problem. Fitting clean vacuum waveforms [48, 49] with these techniques opens the possibility of extracting dozens of QNMs, although their stability and robustness are still not universally defined and depend on the methods used [50]. Additional caveats pertain to the fit’s starting time, which is the point at which perturbation theory is considered a valid description and the BH is described by a Kerr spacetime [51].
The presence of matter is expected in astrophysical environments and could potentially affect our ability to perform precision GW astronomy. A comprehensive examination of this issue has been conducted in Ref. [52] for tractable cases. Because this subject is technically complex, there have been attempts to calculate QNM frequencies from modified/perturbed potentials that are intended to resemble deviations in the underlying perturbation equations [53, 54, 55, 56]. More recently, studies considering an analytical solution for a non-rotating BH that is surrounded by an accretion disk were able to extract QNM frequencies of supermassive BHs, e.g., [57, 58, 59]. A thin shell of matter was recently introduced as a modification to the Regge-Wheeler potential of a Schwarzschild BH in Ref. [60]. They demonstrate how a thin shell is the cause of secondary GW emission and can dramatically shift QNM frequencies for larger mass distributions. A method for phenomenologically inferring accretion disk properties from BNS mergers with features after the ringdown was presented in Ref. [61], and remnant properties have been studied in Ref. [62].
@@
Our work is structured as follows. In Sec. II, we outline the numerical setups used for the gravitational collapse simulations and for the evolution of the BNS mergers. Sec. III introduces the aforementioned ringdown classification scheme applied to the isolated NS cases and shows selected single and binary ringdown simulations. In Sec. IV, we outline our QNM fitting algorithm and follow up with results in Sec. V, for both, isolated and BNS simulations. Our conclusions can be found in Sec. VI. Throughout the paper, we use geometric units and set $G=c=M_{\odot}=1$ (exceptions are indicated throughout).
@@ II.1.1 Isolated stars with differential rotation
The equilibrium models are then computed using the iterative KEH-method (Komatsu-Eriguchi-Hachisu) [75]. The publicly available RNS code [76, 77] implements this scheme and is able to compute equilibrium models of rotating NSs, however, only for stars with uniform rotation. We use a modified version of RNS which can handle differential rotation, using modifications to the KEH-method based on [78, 79].
@@ II.1.2 BNS initial data
For BNS evolutions, we construct initial data using the pseudo-spectral code SGRID [81, 82, 83, 84], which is one of the standard approaches for BNS simulations performed with BAM. SGRID solves the Einstein constraint equations using the extended conformal thin sandwich approach [85, 86]. For the construction of the initial data, we use the piecewise-polytropic fit of [87] of the SLy EOS [88] and add thermal effects through an additional thermal pressure following [89] with $\Gamma_{\rm th}=1.75$.
@@ II.2 Evolution
BAM’s grid infrastructure consists of a variable number of cell-centered nested grids. Each level contains one or more Cartesian boxes with a constant grid spacing and number of points. Due to the 2:1-refinement strategy, the resolution in each finer level is twice the resolution of the parent level. Inner levels can move dynamically according to the technique of ‘moving boxes’ and follow the motion of the NSs in our BNS cases.
@@ III.1 Isolated NS simulations
Our ringdown classification is defined as follows: among the 130 ringdown signals, we observed three distinct signatures. In some cases, we observe a clean ringdown signal that is only cut off once it reaches the numerical noise floor. We define such signals to be of Class I. On the other hand, a signal is of Class II, if we initially find a clean ringdown, which at some point during the ringdown phase gets modified by matter exciting additional oscillations that make the initial ringdown features unrecognizable. Lastly, a signal is of Class III, if it does not exhibit any similarity to a clean ringdown at any time. We show the Weyl scalar $r\Psi_{4}$ for the dominant $(\ell,m)=(2,0)$ mode incorporating the GW information for one representative waveform of each class in Fig. 1. Each model is equipped with an identifier labeling the corresponding rotation parameter and axis ratio. We provide a full list with more details in Appendix A.
@@
We classified all 130 dRNS models through visual inspection (our classification for each model is given in Table LABEL:tab:inittab). In particular, the distinction between Class I and Class II is quite subjective. Our distinguishing criterion between these two classes is that if there were no clear additional oscillations as for the Class II model in Fig. 1, i.e., something that looks like a clean Class I ringdown, then we would look at the amount of peaks present in the modulus of $|r\Psi_{4\ 20}|$. If the model had a suspiciously small number of peaks, then we classified it as a member of Class II, since matter cannot produce high-amplitude oscillations that erase previous ringdown features but rather is of the same amplitude and cuts off these later peaks. In Fig. 2, we show the distribution of classes in the $\hat{A}-r_{p}/r_{e}$ plane. The most models are of Class I (53), followed by Class III (48), and the fewest amount of models are found for Class II (29). Looking back at Fig. 2, we noticed a couple of trends. We find that the smaller the rotation parameter gets, the fewer Class I models we can find, indicating that a larger differential rotation disfavors the presence of a clean ringdown. Especially for models with $\hat{A}\leq 0.2$, we find no Class I models, but all are either of Class II or III, since typically these are the systems with larger rotation and more complex collapse dynamics. Also, we found that models with $r_{p}/r_{e}\geq 0.85$ are exclusively of Class II or Class III. For the Class I models, we identified two “islands” at intermediate and small axis ratios with $\hat{A}>0.2$. However, we will not attempt to draw particular conclusions about possible connections between $\hat{A}$ and $r_{p}/r_{e}$ and our classification, since generally the dynamics of each simulation are complex. Furthermore, we find that, although generally robust and clearly dependent on the presence of matter outside of the formed horizon, the classification for each particular model depends on the exact numerical setup (e.g., resolution, reconstruction, Riemann solver).
@@ III.1.1 Class I
Above the top panel of Fig. 3, we show the time evolution of energy density at different timeslices. We clearly see the oblateness of the star in the xz-plane due to its rotation and axis ratio. At $t-t_{\mathrm{AH}}=-162.0$ ($t_{\mathrm{AH}}$ is the time when an AH is found first) the star sheds some mass in the equatorial plane. Once the gravitational field gets stronger, a BH forms. Shortly afterwards, at $t-t_{\mathrm{AH}}=8.0$, a small asymmetric sphere of matter remains in close proximity to the final BH. We indicate the AH of the BH with a red solid line. At $t-t_{\mathrm{AH}}=58.0$, almost all matter is within the BH. In the last two timeslices, we also show the location of the lightring (white line). We determine the lightring by solving the geodesic equation, shooting geodesics at different starting positions to find a stable orbit. Due to gauge effects, notably the $\Gamma$-driver shift, it grows over time, even though the BH mass stays roughly constant.
Figure 3: The upper row shows the base 10 logarithm of the maximum density on the fifth refinement level in the xy-plane at four different timeslices for the Class I model C8, while the lower row depicts the xz-plane. In the last two panels of each row, the red solid line denotes the AH of the BH, and the white semicircle shows the position of the BH’s lightring. We highlight the corresponding times in the waveform panel of the $(l,m)=(2,0)$ mode, directly below. This panel shows the $(2,0)$ waveform in linear scale (black solid line) and its modulus in log scale (red dotted line) plotted against time with respect to the point of AH formation. Below the $(2,0)$-mode, the $(l,m)=(4,0)$ mode is shown with increased amplitude by a factor of 50. On the last panel, below the $(4,0)$-mode, shown are respectively the baryonic mass on the fifth refinement level outside a spherical approximation of the AH (green) and the mass of the BH (blue).
@@ III.1.2 Class II
In Fig. 4, we see that the star loses some of its material inside the equatorial plane at around $t-t_{\mathrm{AH}}=-135.0$. Once the star collapses, we retain a much heavier disk, which is confined inside the equatorial plane and can reach a baryonic mass of around $\sim 10^{-5}-10^{-6}$. During the ringdown, the dense matter will start to fall into the BH, and at $t-t_{\mathrm{AH}}=195.0$ we only find densities of order $10^{-9}$ around the remnant, which corresponds to a total baryonic mass of $10^{-6}$. However, it is this increased amount of material that leads to additional high-amplitude oscillations in the waveform at around the same time.
@@ III.1.3 Class III
The density evolution is shown above the top panel of Fig. 5. Before the actual collapse, we see some matter outflows in the polar regions at $t-t_{\mathrm{AH}}=205.0$ before the BH forms (we observed a consistent behavior of Class II models showing only equatorial outflows and Class III models showing both equatorial and polar outflows). At $t-t_{\mathrm{AH}}=-96.0$, we find even more massive outflows in the xz-plane, but also additional shedded mass in the xy-plane. Shortly after, the star collapses to a BH. What remains around the remnant are chunks of matter with densities of order $\sim 10^{-9}$, which correspond to a disk mass of $\sim 10^{-6}$. Between the last two snapshots at $t-t_{\mathrm{AH}}=11.0$ and $t-t_{\mathrm{AH}}=90.0$, we can clearly see the backflow of matter towards the BH. This material is constantly exciting the BH.
@@ III.2 BNS simulations
In the following, to apply our classification to more realistic astrophysical scenarios, we investigate three representative BNS simulations and discuss how they fit in the previously defined classification scheme. Due to the increasing computational cost, we cannot easily explore as many simulations as in the isolated NS case.
@@
We start by considering a heavy equal mass BNS system with a total mass of $M_{\text{total}}=3.8M_{\odot}$. In Fig. 6, we show a similar plot as in Sec. III. Due to the change in symmetry (isolated to binary), we now report the dominant (2,2)-mode of $r\Psi_{4}$ (extracted at $r=400$), and the baryonic mass outside of the AH, as well as the BH mass. Due to the high mass of the system, no hypermassive NS is formed, but the system undergoes a prompt collapse to a BH. A snapshot of the BH during the ringdown is shown at $t-t_{\mathrm{AH}}=191.0$. In this case, only small amounts of baryonic mass of the order of $10^{-5}$ remain outside. However, we still see that matter falls into the BH for around half of the ringdown signal which results in the BH mass increasing by around 16% in the early stages after formation. Because of this considerable change in mass, it is remarkable that the ringdown signal is very clean and long. The ringdown signal itself extends over a range of $\sim 300$ after BH formation and shows 14-15 clear ringdown peaks in the modulus. Relating back to the previous classification scheme, this corresponds to a Class I ringdown.
@@
The simulation is shown in Fig. 7. Due to the different component masses, the heavier NS disrupts the lighter one earlier, as seen in the snapshot at $t-t_{\mathrm{AH}}=-68.0$. In our simulation (not visible here), the stars merge at a bit earlier than in the previous case, but BH formation is slightly delayed. It is interesting that the ringdown does not have a typical envelope, but rather some peaks that are slightly lifted up or down with respect to the others. We suspect that it is due to the heavy disk with baryonic mass of around $10^{-2}$ surrounding the remnant. The BH mass grows around 13% during the ringdown phase, which might introduce additional slight perturbations of the BH, which lead to lifted peaks. Although the ringdown looks like Class I, some peaks look odd and a possible Class II ringdown could also be a valid description due to the large baryonic mass outside. The QNM extraction is, however, still possible and the results are reported in Sec. V.5.
@@
As our last example, we consider an equal mass binary with a lower total mass of $M_{\text{tot}}=2.7M_{\odot}$. The waveform is extracted at $r=300$ on the 2nd refinement level, as for example, the resolution on the 0th refinement level would not have been sufficient to resolve the ringdown.
Looking at the density evolution in Fig. 8, we find that at around $t-t_{\mathrm{AH}}\approx-2300$, both NSs come into contact shortly before the merger. However, due to the low mass of the individual NSs, we do not observe a prompt collapse to a BH, but rather a phase where a hypermassive NS [106, 107, 108] forms, which survives for $\sim 2300$, until it collapses. The mass outside the AH is around $10^{-2}$, and the BH mass only increases by about 8% in the range of $\sim 100$ after its formation. In the waveform, we see that the modulus of $r\Psi_{4\ 22}$ in log-scale does not show a lot of ringdown peaks and the ringdown is rather short compared to the two previous cases. For that reason we rather classify it as Class II since matter seems to suppress the length of the signal significantly.

### v2

@@ Abstract
Black hole spectroscopy is an important pillar when studying gravitational waves from black holes and enables tests of general relativity. Most of the gravitational wave signals observed over the last decade originate from binary black hole systems. Binary neutron star or black hole neutron-star systems are rarer but of particular interest for the next-generation ground-based gravitational wave detectors. These events offer the exciting possibility of studying matter effects on the ringdown of “dirty black holes.” In this work, we ask the question; Does matter matter? Using numerical-relativity, we simulate a wide range of collapsing neutron stars producing matter environments, both in isolated scenarios and in binary mergers. Qualitatively, the resulting ringdown signals can be classified into “clean,” “modified,” and “distorted” cases, depending on the amount of matter that is present. We apply standard strategies for extracting quasinormal modes of clean signals, using both theory-agnostic and theory-specific assumptions. Even in the presence of matter, possible modifications of quasinormal modes seem to be dominated by ringdown modeling systematics. We find that incorporating multiple quasinormal modes allows one to drastically reduce mismatches and errors in estimating the final black hole mass at early times. If not treated carefully, deviations in the fundamental quasinormal mode might artificially be overestimated and falsely attributed to the presence of matter or violations of general relativity.
@@ I Introduction
By concentrating on the ringdown, it is possible to test the no-hair theorem’s assumptions. A BH formed from the dynamical processes involving compact objects “rings down” from an initial state of nonequilibrium and emits GW radiation in form of damped sinusoids. Pioneering works first targeted nonrotating BHs [24, 25, 26, 27], and then rotating ones [28]. They discovered how a BH responds to an incoming GW pulse: it oscillates with characteristic damped frequencies. The damping is due to GWs in the vicinity of the horizon falling into the BH, while GWs far away leave the system. The ringdown is a consequence of both conditions. This stage is anticipated to occur as a result of the coalescence of compact objects or in other instances where BH formation occurs, such as the collapse of a perturbed NS. Seminal studies probing the nonlinear BH ringdown with numerical relativity can be found in Ref. [29, 30, 31].
@@
In vacuum and in GR, the QNM spectrum is well known [34], while it requires strong simplifications to examine the QNM content in matter environments to make the equations decouple and make the analysis tractable. Maintaining the vacuum frequencies at a constant value from the start, allows for a simplified extraction of the complex-valued amplitudes $C_{\ell mn}$ from the signal. The amplitudes are fitted using standard least-squares fitting procedures [43, 44, 45], while the variable projection method [46, 47] is employed for the nonlinear fitting problem. Fitting clean vacuum waveforms [48, 49] with these techniques opens the possibility of extracting dozens of QNMs, although their stability and robustness are still not universally defined and depend on the methods used [50]. Additional caveats pertain to the fit’s starting time, which is the point at which perturbation theory is considered a valid description and the BH is described by a Kerr spacetime [51].
The presence of matter is expected in astrophysical environments and could potentially affect our ability to perform precision GW astronomy. A comprehensive examination of this issue has been conducted in Ref. [52] for tractable cases. Because this subject is technically complex, there have been attempts to calculate QNM frequencies from modified/perturbed potentials that are intended to resemble deviations in the underlying perturbation equations [53, 54, 55, 56]. More recently, studies considering an analytical solution for a nonrotating BH that is surrounded by an accretion disk were able to extract QNM frequencies of supermassive BHs, e.g., [57, 58, 59]. A thin shell of matter was recently introduced as a modification to the Regge-Wheeler potential of a Schwarzschild BH in Ref. [60]. They demonstrate how a thin shell is the cause of secondary GW emission and can dramatically shift QNM frequencies for larger mass distributions. A method for phenomenologically inferring accretion disk properties from BNS mergers with features after the ringdown was presented in Ref. [61], and remnant properties have been studied in Ref. [62]. Additionally, Ref. [63] used BNS merger simulations to attest whether QNM amplitudes can aid in determining pre-merger binary properties such as the mass ratio.
@@
Our work is structured as follows. In Sec. II, we outline the numerical setups used for the gravitational collapse simulations and for the evolution of the BNS mergers. Section III introduces the aforementioned ringdown classification scheme applied to the isolated NS cases and shows selected single and binary ringdown simulations. In Sec. IV, we outline our QNM fitting algorithm and follow up with results in Sec. V, for both, isolated and BNS simulations. Our conclusions can be found in Sec. VI. Throughout the paper, we use geometric units and set $G=c=M_{\odot}=1$ (exceptions are indicated throughout).
@@ II.1.1 Isolated stars with differential rotation
The equilibrium models are then computed using the iterative Komatsu-Eriguchi-Hachisu (KEH) method [76]. The publicly available RNS code [77, 78] implements this scheme and is able to compute equilibrium models of rotating NSs, however, only for stars with uniform rotation. We use a modified version of RNS which can handle differential rotation, using modifications to the KEH method based on [79, 80].
@@ II.1.2 BNS initial data
For BNS evolutions, we construct initial data using the pseudospectral code SGRID [82, 83, 84, 85], which is one of the standard approaches for BNS simulations performed with BAM. SGRID solves the Einstein constraint equations using the extended conformal thin sandwich approach [86, 87]. For the construction of the initial data, we use the piecewise-polytropic fit of [88] of the SLy EOS [89] and add thermal effects through an additional thermal pressure following [90] with $\Gamma_{\rm th}=1.75$.
@@ II.2 Evolution
BAM’s grid infrastructure consists of a variable number of cell-centered nested grids. Each level contains one or more Cartesian boxes with a constant grid spacing and number of points. Due to the 2:1 refinement strategy, the resolution in each finer level is twice the resolution of the parent level. Inner levels can move dynamically according to the technique of ‘moving boxes’ and follow the motion of the NSs in our BNS cases.
@@ III.1 Isolated NS simulations
Our ringdown classification is defined as follows. Among the 130 ringdown signals, we observed three distinct signatures. In some cases, we observe a clean ringdown signal that is only cut off once it reaches the numerical noise floor. We define such signals to be of Class I. On the other hand, a signal is of Class II, if we initially find a clean ringdown, which at some point during the ringdown phase gets modified by matter exciting additional oscillations that make the initial ringdown features unrecognizable. Lastly, a signal is of Class III, if it does not exhibit any similarity to a clean ringdown at any time. We show the Weyl scalar $r\Psi_{4}$ for the dominant $(\ell,m)=(2,0)$-mode incorporating the GW information for one representative waveform of each class in Fig. 1. Each model is equipped with an identifier labeling the corresponding rotation parameter and axis ratio. We provide a full list with more details in Appendix A.
@@
We classified all 130 dRNS models through visual inspection (our classification for each model is given in Table LABEL:tab:inittab). In particular, the distinction between Class I and Class II is quite subjective. Our distinguishing criterion between these two classes is that if there were no clear additional oscillations as for the Class II model in Fig. 1, i.e., something that looks like a clean Class I ringdown, then we would look at the amount of peaks present in the modulus of $|r\Psi_{4\ 20}|$. If the model had a suspiciously small number of peaks, then we classified it as a member of Class II, since matter cannot produce high-amplitude oscillations that erase previous ringdown features but rather is of the same amplitude and cuts off these later peaks. In Fig. 2, we show the distribution of classes in the $\hat{A}-r_{p}/r_{e}$ plane. The most models are of Class I (53), followed by Class III (48), and the fewest amount of models are found for Class II (29). Looking back at Fig. 2, we noticed a couple of trends. We find that the smaller the rotation parameter gets, the fewer Class I models we can find, indicating that a larger differential rotation disfavors the presence of a clean ringdown. Especially for models with $\hat{A}\leq 0.2$, we find no Class I models, but all are either of Class II or III, since typically these are the systems with larger rotation and more complex collapse dynamics. Also, we found that models with $r_{p}/r_{e}\geq 0.85$ are exclusively of Class II or Class III. For the Class I models, we identified two “islands” at intermediate and small axis ratios with $\hat{A}>0.2$. However, we will not attempt to draw particular conclusions about possible connections between $\hat{A}$ and $r_{p}/r_{e}$ and our classification, since generally the dynamics of each simulation are complex.
Furthermore, we find that, although generally robust and clearly dependent on the presence of matter outside of the formed horizon, the classification for each particular model depends on the exact numerical setup (e.g., resolution, reconstruction, Riemann solver). [^2] To be more explicit, for the 130 isolated NS simulations in this work, we first performed respective simulations at a lower resolution. We found that for about half of the models a different collapse dynamics compared to the high-resolution runs was present. This in turn influenced the ringdown morphology and frequency extraction. The relatively large difference between the different resolutions comes from the fact that the employed setups probe the collapse dynamics near the critical point of stability and is introduced through a small perturbation. Larger perturbations would reduce the influence of the resolution, but would also introduce larger spurious constraint violations. Most importantly, the obtained classification of the ringdown signal is directly linked to the outflow of matter independent of the resolution, i.e., setups for which a large amount of matter is released before merger (independent of the resolution) will lead to Class III evolutions, while setups with a small amount of mass ejection will lead to Class I setups. Furthermore, we find that different reconstruction schemes and Riemann solvers seemed to have only a negligible influence on the results. The BNS cases on the other hand were robust, and the numerical setup only influenced the time of merger, but not the classification.
@@ III.1.1 Class I
Above the top panel of Fig. 3, we show the time evolution of energy density at different time slices. We clearly see the oblateness of the star in the $xz$-plane due to its rotation and axis ratio. At $t-t_{\mathrm{AH}}=-162.0$ ($t_{\mathrm{AH}}$ is the time when an AH is found first) the star sheds some mass in the equatorial plane. Once the gravitational field gets stronger, a BH with mass $M_{\mathrm{BH}}=1.861M_{\odot}$ and spin $a_{\mathrm{BH}}=0.519$ forms. Shortly afterwards, at $t-t_{\mathrm{AH}}=8.0$, a small asymmetric sphere of matter remains in close proximity to the final BH. We indicate the AH of the BH with a red solid line. At $t-t_{\mathrm{AH}}=58.0$, almost all matter is within the BH. In the last two time slices, we also show the location of the lightring (white line). We determine the lightring by solving the geodesic equation, shooting geodesics at different starting positions to find a stable orbit. Due to gauge effects, notably the $\Gamma$-driver shift, it grows over time, even though the BH mass stays roughly constant.
Figure 3: The upper row shows the base 10 logarithm of the maximum density on the fifth refinement level in the $xy$-plane at four different time slices for the Class I model C8, while the lower row depicts the $xz$-plane. In the last two panels of each row, the red solid line denotes the AH of the BH, and the white semicircle shows the position of the BH’s lightring. We highlight the corresponding times in the waveform panel of the $(l,m)=(2,0)$-mode, directly below. This panel shows the $(2,0)$ waveform in linear scale (black solid line) and its modulus in log scale (red dotted line) plotted against time with respect to the point of AH formation. Below the $(2,0)$-mode, the $(l,m)=(4,0)$-mode is shown with increased amplitude by a factor of 50. On the last panel, below the $(4,0)$-mode, shown are respectively the baryonic mass on the fifth refinement level outside a spherical approximation of the AH (green) and the mass of the BH (blue).
@@ III.1.2 Class II
In Fig. 4, we see that the star loses some of its material inside the equatorial plane at around $t-t_{\mathrm{AH}}=-135.0$. Once the star collapses, we retain a much heavier disk, which is confined inside the equatorial plane and can reach a baryonic mass of around $\sim 10^{-5}-10^{-6}$. The BH reaches a mass of $M_{\mathrm{BH}}=2.104M_{\odot}$ and a spin of $a_{\mathrm{BH}}=0.684$, which is respectively larger than for the C8 model. During the ringdown, the dense matter will start to fall into the BH, and at $t-t_{\mathrm{AH}}=195.0$ we only find densities of order $10^{-9}$ around the remnant, which corresponds to a total baryonic mass of $10^{-6}$. However, it is this increased amount of material that leads to additional high-amplitude oscillations in the waveform at around the same time.
@@ III.1.3 Class III
The density evolution is shown above the top panel of Fig. 5. Before the actual collapse, we see some matter outflows in the polar regions at $t-t_{\mathrm{AH}}=205.0$ before the BH forms (we observed a consistent behavior of Class II models showing only equatorial outflows and Class III models showing both equatorial and polar outflows). At $t-t_{\mathrm{AH}}=-96.0$, we find even more massive outflows in the $xz$-plane, but also additional shedded mass in the $xy$-plane. Shortly after, the star collapses to a BH with mass $M_{\mathrm{BH}}=1.759M_{\odot}$ and spin $a_{\mathrm{BH}}=0.336$. What remains around the remnant are chunks of matter with densities of order $\sim 10^{-9}$, which correspond to a disk mass of $\sim 10^{-6}$. Between the last two snapshots at $t-t_{\mathrm{AH}}=11.0$ and $t-t_{\mathrm{AH}}=90.0$, we can clearly see the backflow of matter towards the BH. This material is constantly exciting the BH.
@@ III.2 BNS simulations
In the following, to apply our classification to more realistic astrophysical scenarios, we investigate three representative BNS simulations and discuss how they fit in the previously defined classification scheme. Studies of the BH ringdown in BNS simulations have been presented in, e.g.,  [105, 63]. Studies of BHNS systems can be found in [106] and are in agreement with the findings of [107].
@@
We start by considering a heavy equal mass BNS system with a total mass of $M_{\text{total}}=3.8M_{\odot}$. In Fig. 6, we show a similar plot as in Sec. III. Due to the change in symmetry (isolated to binary), we now report the dominant (2,2)-mode of $r\Psi_{4}$ (extracted at $r=400$), and the baryonic mass outside of the AH, as well as the BH mass. Due to the high mass of the system, no hypermassive NS is formed, but the system undergoes a prompt collapse to a BH with a mass of $M_{\mathrm{BH}}=2.958M_{\odot}$ after formation, which increases shortly after reaches a plateau of $M_{\mathrm{BH}}=3.659M_{\odot}$. Likewise, the spin increases from initially $a_{\mathrm{BH}}=0.628$ after merger to $a_{\mathrm{BH}}=0.726$. A snapshot of the BH during the ringdown is shown at $t-t_{\mathrm{AH}}=191.0$. In this case, only small amounts of baryonic mass of the order of $10^{-5}$ remain outside. However, we still see that matter falls into the BH for around half of the ringdown signal which results in the BH mass increasing by around 16% in the early stages after formation. Because of this considerable change in mass, it is remarkable that the ringdown signal is very clean and long. The ringdown signal itself extends over a range of $\sim 300$ after BH formation and shows 14-15 clear ringdown peaks in the modulus. Relating back to the previous classification scheme, this corresponds to a Class I ringdown.
@@
The simulation is shown in Fig. 7. Due to the different component masses, the heavier NS disrupts the lighter one earlier, as seen in the snapshot at $t-t_{\mathrm{AH}}=-68.0$. In our simulation (not visible here), the stars merge at a bit earlier than in the previous case, but BH formation is slightly delayed. After accreting some of the surrounding matter, the BH reaches a mass of $M_{\mathrm{BH}}=3.116M_{\odot}$ and a spin of $a_{\mathrm{BH}}=0.778$. It is interesting that the ringdown does not have a typical envelope, but rather some peaks that are slightly lifted up or down with respect to the others. We suspect that it is due to the heavy disk with baryonic mass of around $10^{-2}$ surrounding the remnant. The BH mass grows around 13% during the ringdown phase, which might introduce additional slight perturbations of the BH, which lead to lifted peaks. Although the ringdown looks like Class I, some peaks look odd and a possible Class II ringdown could also be a valid description due to the large baryonic mass outside. The QNM extraction is, however, still possible and the results are reported in Sec. V.5.
@@
As our last example, we consider an equal mass binary with a lower total mass of $M_{\text{tot}}=2.7M_{\odot}$. The waveform is extracted at $r=300$ on the second refinement level, as for example, the resolution on the zeroth refinement level would not have been sufficient to resolve the ringdown.
Looking at the density evolution in Fig. 8, we find that at around $t-t_{\mathrm{AH}}\approx-2300$, both NSs come into contact shortly before the merger. However, due to the low mass of the individual NSs, we do not observe a prompt collapse to a BH, but rather a phase where a hypermassive NS [110, 111, 112] forms, which survives for $\sim 2300$, until it collapses. The mass outside the AH is around $10^{-2}$, and the BH mass only increases by about 8% in the range of $\sim 100$ after its formation. After the BH settles into equilibrium, its mass reached $M_{\mathrm{BH}}=2.522M_{\odot}$ and its spin $a_{\mathrm{BH}}=0.661$. In the waveform, we see that the modulus of $r\Psi_{4\ 22}$ in log-scale does not show a lot of ringdown peaks and the ringdown is rather short compared to the two previous cases. For that reason we rather classify it as Class II since matter seems to suppress the length of the signal significantly.